# AtomR: Atomic Operator-Empowered Large Language Models for Heterogeneous Knowledge Reasoning

### v3

@@ Abstract.
Despite the outstanding capabilities of large language models (LLMs), knowledge-intensive reasoning still remains a challenging task due to LLMs’ limitations in compositional reasoning and the hallucination problem. A prevalent solution is to employ chain-of-thought (CoT) with retrieval-augmented generation (RAG), which first formulates a reasoning plan by decomposing complex questions into simpler sub-questions, and then applies iterative RAG at each sub-question. However, prior works exhibit two crucial problems: inadequate reasoning planning and poor incorporation of heterogeneous knowledge. In this paper, we introduce AtomR, a framework for LLMs to conduct accurate heterogeneous knowledge reasoning at the atomic level. Inspired by how knowledge graph query languages model compositional reasoning through combining predefined operations, we propose three atomic knowledge operators, a unified set of operators for LLMs to retrieve and manipulate knowledge from heterogeneous sources. First, in the reasoning planning stage, AtomR decomposes a complex question into a reasoning tree where each leaf node corresponds to an atomic knowledge operator, achieving question decomposition that is highly fine-grained and orthogonal. Subsequently, in the reasoning execution stage, AtomR executes each atomic knowledge operator, which flexibly selects, retrieves, and operates atomic level knowledge from heterogeneous sources. We also introduce BlendQA, a challenging benchmark specially tailored for heterogeneous knowledge reasoning. Experiments on three single-source and two multi-source datasets show that AtomR outperforms state-of-the-art baselines by a large margin, with F1 score improvements of $9.4\%$ on 2WikiMultihop and $9.5\%$ on BlendQA. We release our code and datasets [^1] https://github.com/THU-KEG/AtomR.git.
@@
copyright: acmlicensedjournalyear: 2024conference: ; Under Review; THUKEGccs: Information systems Question answeringccs: Computing methodologies Natural language generation
@@ 1. Introduction
Knowledge-intensive reasoning is a challenging task that requires the ability to perform compositional reasoning over vast volumes of knowledge, demanding skills such as multi-hop inference, comparison and calculation (Yin et al., 2022; Gu et al., 2021; Cao et al., 2022; Trivedi et al., 2022). Although recent advancements of Large Language Models (LLMs) (OpenAI, 2023) have enabled them to excel in various natural language processing tasks(Zhao et al., 2023; Bang et al., 2023), it is still arduous for LLMs to perform knowledge-intensive reasoning due to their weaknesses in implicit compositional reasoning (Press et al., 2023; Dziri et al., 2024) and the hallucination problem (Huang et al., 2023; Bang et al., 2023). To address these issues, recent works (Trivedi et al., 2023; Press et al., 2023; Xu et al., 2024; Cao et al., 2023) propose to leverage chain-of-thought (CoT) (Wei et al., 2022) reasoning with retrieval augmented generation (RAG): specifically, to first perform question decomposition on complex questions into simpler sub-questions, thereby formulating an explicit reasoning plan, and then retrieve external knowledge for sub-question answering, thereby alleviating LLM hallucination.
@@
However, there are still three main challenges. $\mathbf{\mathcal{C}1}$. Sub-optimal reasoning planning due to inadequate question decomposition. Existing works have explored question decomposition into different structures, such as chain (Trivedi et al., 2023; Press et al., 2023; Shao et al., 2023; Xu et al., 2024) and tree structures (Yao et al., 2023b; Cao et al., 2023). However, despite constraining the decomposition structure, previous works leverage LLMs to generate sub-questions freely, imposing no supervision over the atomicity or semantic function of sub-questions. This results in the generation of coarse-grained sub-questions as well as sub-questions with inaccurate functionalities, leading to sub-optimal reasoning paths and mistakes. $\mathbf{\mathcal{C}2}$. The limited support for multiple heterogeneous knowledge sources. Online web pages, local text corpora, and structured knowledge bases contain rich knowledge that complement each other. However, most existing works only retrieve from a fixed knowledge source (Trivedi et al., 2023; Xu et al., 2024). A few (Li et al., 2024) support multiple knowledge sources, but fall short in cross-source retrieval at the step level, as well as effective knowledge manipulation. $\mathbf{\mathcal{C}3}$. The absence of high quality datasets built on heterogeneous knowledge sources. Existing benchmarks that are built on multiple knowledge sources either have a narrow knowledge scope, typically only encompassing encyclopedic knowledge from Wikipedia and Wikidata (Zhao et al., 2024; Zhang et al., 2024; Yao et al., 2023a), or lack instance-level design, allowing individual questions to be answered using a single knowledge source without requiring cross-knowledge source querying and aggregation (Christmann et al., 2024; Yang et al., 2024).
@@ 2. Related Work
Knowledge-intensive reasoning requires both massive world knowledge and compositional reasoning skills. It reflects a model’s ability of using extensive knowledge to solve real-world problems: more specifically, to deduce accurate conclusions by effectively aggregating scattered pieces of knowledge. Knowledge-intensive reasoning can be evaluated through a wide range of tasks, such as multi-hop question answering (QA) (Yang et al., 2018; Ho et al., 2020; Trivedi et al., 2022), fact verification (Thorne et al., 2018; Wadden et al., 2020), and knowledgeable open dialogue (Dinan et al., 2019; Yu et al., 2022). In this work, we focus on the multi-hop QA task for its effective modeling of complex reasoning structures. Furthermore, the fact that world knowledge is dispersed across diverse structures and sources calls for knowledge-intensive reasoning systems to incorporate multiple heterogeneous knowledge sources. Some recent works (Li et al., 2024; Lehmann et al., 2024) explore the integration of heterogeneous knowledge retrieval. However, these works retrieve knowledge from the same static set of sources throughout each entire question, instead of dynamically retrieving from different suitable source(s) at each reasoning step, which is common in real-world scenarios. We also observe that existing multi-source benchmarks (Christmann et al., 2024; Yang et al., 2024) fail to assess such step-level cross-knowledge source retrieval. Hence, we design AtomR to support flexible knowledge source selection at each atomic operator, and construct BlendQA to evaluate cross-knowledge source reasoning.
@@
Retrieval-augmented LLMs have shown promising potential in knowledge-intensive reasoning. Early works adopt a one-time retrieval strategy (Izacard et al., 2023; Borgeaud et al., 2022), where the retriever is only called once to solve each question. This is often insufficient for solving multi-hop questions, which requires retrieval at each reasoning step. Hence, more recent approaches turn to multi-time retrieval strategies. IRCoT (Trivedi et al., 2023) and ITER-RETGEN (Shao et al., 2023) iteratively calls retrieval at each CoT step based on the previous steps’ results. Self-Ask (Press et al., 2023) iteratively generates a chain of sub-questions and performs retrieval at each sub-question. ProbTree (Cao et al., 2023) proposes to first formulate a global question decomposition plan in tree structure, then recursively perform retrieval at each sub-question. However, existing multi-time retrieval frameworks leverage LLMs to generate sub-questions in a free-form manner, imposing no supervision over each sub-question’s atomicity and functionality. Compared to previous methods, AtomR uses the combination of three predefined atomic knowledge operators to model complex reasoning procedures, achieving question decomposition that is highly fine-grained and orthogonal.
@@ 3. Methodology
In this section, we introduce the design of AtomR. The overall architecture of AtomR is illustrated in Figure 2. At the core of our framework are three fundamental Atomic Knowledge Operators (Section 3.1), which we design to retrieve and manipulate knowledge at the atomic level. These operators then steer the process of AtomR through two main stages: (1) Atomic Reasoning Planning (Section 3.2), where the system decomposes the input question into a fine-grained Atomic Reasoning Tree (ART), and (2) Atomic Reasoning Execution (Section 3.3), where the system performs bottom-up reasoning across multiple heterogeneous knowledge sources at each atomic node. We first introduce the design of our atomic knowledge operators in Section 3.1, then detail the two main stages of our framework in Sections 3.2 and 3.3. The LLM prompts of AtomR are included in Appendix F. The benchmark construction of BlendQA will be introduced later in Section 4.1.2.
@@ 3.1. Atomic Knowledge Operators
Compositional reasoning, the ability to construct complex reasoning procedures by combining simple components, is a fundamental property of human thought (Fodor and Pylyshyn, 1988; Fodor, 2002). Knowledge graphs (KGs) effectively model the compositionality of knowledge, breaking down complex world knowledge into combinations of atomic elements: entities, relations, and attributes. Furthermore, KG query languages use predefined operations to manipulate the atomic elements in KGs. In this light, we ask the question: Can we use a set of predefined operators to atomically manipulate knowledge and enhance LLM compositional reasoning?
We investigate the essential operations of existing KG query languages, and observe that all operations can be distilled into combinations of three basic atomic knowledge operators: Search, Relate, and Filter. Table 1 illustrates how operations of three widely-used KG query languages, SPARQL, Cypher, and KoPL (Pérez et al., 2006; Francis et al., 2018; Cao et al., 2022), can be induced into AtomR’s three atomic knowledge operators. Each operator corresponds to a unique, orthogonal function. We implement each operator on multiple heterogeneous knowledge sources with an adaptive LLM executor, as shown in Figure 2(2)(iii).
@@
During execution, similar as Search, Relate first initiates multi-source knowledge retrieval using the concatenated parameters as the query, then inputs the retrieved knowledge and corresponding sub-question into the adaptive LLM executor with an in-context learning prompt (Figure 8). Finally, the LLM executor outputs the answer: either an entity list for case (1), an attribute value for case (2), or a relation for case (3).
@@
The overlapping coefficient of each entity is then compared to a hyperparameter threshold $t$. Entities with $O_{i}<t$ will be directly discarded to remove excessive noise from the retrieved knowledge. Then, the remaining entities and passages are inputted to the adaptive LLM executor to perform entity filtering via in-context learning (Figure 9). Finally, the LLM executor outputs the filtered entity list.
@@ 3.3.1. Leaf Node Reasoning
Finally, with the retrieved knowledge, AtomR executes the atomic knowledge operator. The detailed functionality and procedures of each atomic knowledge operator has been explained in Section 3.1. To ensure flexibility and robustness, we utilize an Adaptive LLM Executor, instead of static symbolic code, for atomic operator execution. Specifically, we employ in-context learning to prompt an LLM to perform the desired knowledge operation (Figures 7 to  9). Finally, the output of the adaptive LLM executor serves as the local answer for the atomic operator as well as its associated sub-question.
@@ 3.3.2. Parent Node Reasoning
$a_{i}=ChildAnswerReasoning(q_{i},[q_{c_{1}},a_{c_{1}},q_{c_{2}},a_{c_{2}},...])$
where $q_{i}$ is the sub-question for the current node and $[q_{c_{1}},a_{c_{1}},...]$ is the list of question-answer pairs for $q_{i}$’s child nodes. For example, in Figure 2(1), sub-question 1 can be answered by synthesizing the answers of its child nodes 4 and 5. Child answer reasoning is achieved through in-context learning, where the LLM is provided with the child question-answer pairs along with an instruction prompt (Figure 12).
@@
$a_{i}=SiblingAnswerReasoning(q_{i},[q_{s_{1}},a_{s_{1}},q_{s_{2}},a_{s_{2}},...])$
where $[q_{s_{1}},a_{s_{1}},...]$ is the list of question-answer pairs for $q_{i}$’s sibling nodes that are referenced in the current question $q_{i}$. For example, in Figure 2(1), sub-question 3 can be answered by analyzing answers of its referenced sibling node 2. Sibling answer reasoning is also achieved through LLM in-context learning (Figure 12).
@@
Each parent node is primarily resolved through either child answer reasoning or sibling answer reasoning. However, if a parent node fails to obtain an answer, direct RAG reasoning is employed. Parallel to leaf node reasoning, direct rag reasoning first performs dynamic knowledge source selection for the current sub-question $q_{i}$, then initiates multi-source knowledge retrieval with $q_{i}$ as the query. Finally, given $q_{i}$ and the retrieved knowledge, an LLM is employed to conduct standard RAG through in context learning (Figure 13). This design ensures that while AtomR primarily focuses on knowledge retrieval at the leaf nodes, it can also flexibly access external knowledge at inner nodes when necessary.
@@ 4. Experiments
Table 2. Results for Single-Source Reasoning. Token-level F1 is reported, with the best results in bold and the second best results underlined. ()($\uparrow$)( ↑ ) indicates the overall gain of AtomR compared to the second-best baseline for each dataset.
@@
AtomR (ours) 71.27$(\uparrow 5.4)$ 68.96 82.07 78.72$(\uparrow 9.4)$ 59.07 80.04 97.48 93.33 36.11$(\uparrow 1.2)$ 45.63 29.94 20.15
@@
AtomR (ours) 43.32$(\uparrow 9.5)$ 45.63 39.69 45.22 71.39$(\uparrow 6.6)$ 70.10 68.51 81.71 64.95

### v4

@@ Abstract.
Despite the outstanding capabilities of large language models (LLMs), knowledge-intensive reasoning still remains a challenging task due to LLMs’ limitations in compositional reasoning and the hallucination problem. A prevalent solution is to employ chain-of-thought (CoT) with retrieval-augmented generation (RAG), which first formulates a reasoning plan by decomposing complex questions into simpler sub-questions, and then applies iterative RAG at each sub-question. However, prior works exhibit two crucial problems: inadequate reasoning planning and poor incorporation of heterogeneous knowledge. In this paper, we introduce AtomR, a framework for LLMs to conduct accurate heterogeneous knowledge reasoning at the atomic level. Inspired by how knowledge graph query languages model compositional reasoning through combining predefined operations, we propose three atomic knowledge operators, a unified set of operators for LLMs to retrieve and manipulate knowledge from heterogeneous sources. First, in the reasoning planning stage, AtomR decomposes a complex question into a reasoning tree where each leaf node corresponds to an atomic knowledge operator, achieving question decomposition that is highly fine-grained and orthogonal. Subsequently, in the reasoning execution stage, AtomR executes each atomic knowledge operator, which flexibly selects, retrieves, and operates atomic level knowledge from heterogeneous sources. We also introduce BlendQA, a challenging benchmark specially tailored for heterogeneous knowledge reasoning. Experiments on three single-source and two multi-source datasets show that AtomR outperforms state-of-the-art baselines by a large margin, with absolute F1 score improvements of $9.4\%$ on 2WikiMultihop and $9.5\%$ on BlendQA. We release our code and data [^1] https://github.com/THU-KEG/AtomR.git.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736849isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Question answeringccs: Computing methodologies Natural language generation
KDD Availability Link:
The source code and data of this paper has been made publicly available at %****␣main.tex␣Line␣300␣****%Artifact␣Badgehttps://doi.org/10.5281/zenodo.15528472.
@@ 1. Introduction
Knowledge-intensive reasoning is a challenging task that requires the ability to perform compositional reasoning over vast volumes of knowledge, demanding skills such as multi-hop inference, comparison and calculation (Yang et al., 2024a; Gu et al., 2021; Cao et al., 2022; Trivedi et al., 2022). Although recent advancements of Large Language Models (LLMs) (Achiam et al., 2023) have enabled them to excel in various natural language processing tasks(Zhao et al., 2023; Bang et al., 2023), it is still arduous for LLMs to perform knowledge-intensive reasoning due to their weaknesses in implicit compositional reasoning (Press et al., 2023; Dziri et al., 2023) and the hallucination problem (Izacard et al., 2023a; Bang et al., 2023). To address these issues, recent works (Trivedi et al., 2023; Press et al., 2023; Xu et al., 2024; Cao et al., 2023) propose to leverage chain-of-thought (CoT) (Wei et al., 2022) reasoning with retrieval augmented generation (RAG): specifically, to first perform question decomposition on complex questions into simpler sub-questions, thereby formulating an explicit reasoning plan, and then retrieve external knowledge for sub-question answering, thereby alleviating LLM hallucination.
@@
However, there are still three main challenges. $\mathbf{\mathcal{C}1}$. Sub-optimal reasoning planning due to inadequate question decomposition. Existing works have explored question decomposition into different structures, such as chain (Trivedi et al., 2023; Press et al., 2023; Shao et al., 2023; Xu et al., 2024) and tree structures (Yao et al., 2023b; Cao et al., 2023). However, despite constraining the decomposition structure, previous works leverage LLMs to generate sub-questions freely, imposing no supervision over the atomicity or semantic function of sub-questions. This results in the generation of coarse-grained sub-questions as well as sub-questions with inaccurate functionalities, leading to sub-optimal reasoning paths and mistakes. $\mathbf{\mathcal{C}2}$. The limited support for multiple heterogeneous knowledge sources. Heterogeneous knowledge sources such as online web pages, local text corpora, and structured knowledge bases contain rich knowledge that complement each other. For example, to resolve the complex question “How many studio albums has Shakira released between 2000 and 2010?”, an effective solution is to first determine who “Shakira” is using a web search engine, then retrieve the albums released by “Shakira” from a structured knowledge base. However, most existing frameworks only retrieve from a fixed knowledge source (Trivedi et al., 2023; Xu et al., 2024). A few (Li et al., 2024) support multiple knowledge sources, but fall short in cross-source retrieval at the step level, as well as effective knowledge manipulation. $\mathbf{\mathcal{C}3}$. The absence of high quality datasets built on heterogeneous knowledge sources. Existing benchmarks that are built on multiple knowledge sources either have a narrow knowledge scope, typically only encompassing encyclopedic knowledge from Wikipedia and Wikidata (Zhao et al., 2024; Zhang et al., 2024; Yao et al., 2023a), or lack instance-level design, allowing individual questions to be answered using a single knowledge source without requiring cross-knowledge source querying and aggregation (Christmann et al., 2024; Yang et al., 2024b).
@@ 2. Related Work
Knowledge-intensive reasoning requires both massive world knowledge and compositional reasoning skills. It reflects a model’s ability of using extensive knowledge to solve real-world problems: more specifically, to deduce accurate conclusions by effectively aggregating scattered pieces of knowledge. Knowledge-intensive reasoning can be evaluated through a wide range of tasks, such as multi-hop question answering (QA) (Yang et al., 2018; Ho et al., 2020; Trivedi et al., 2022), fact verification (Thorne et al., 2018; Wadden et al., 2020), and knowledgeable open dialogue (Dinan et al., 2019; Yu et al., 2022). In this work, we focus on the multi-hop QA task for its effective modeling of complex reasoning structures. Furthermore, the fact that world knowledge is dispersed across diverse structures and sources calls for knowledge-intensive reasoning systems to incorporate multiple heterogeneous knowledge sources. Some recent works (Li et al., 2024; Lehmann et al., 2024) explore the integration of heterogeneous knowledge retrieval. However, these works retrieve knowledge from the same static set of sources throughout each entire question, instead of dynamically retrieving from different suitable source(s) at each reasoning step, which is common in real-world scenarios. We also observe that existing multi-source benchmarks (Christmann et al., 2024; Yang et al., 2024b) fail to assess such step-level cross-knowledge source retrieval. Hence, we design AtomR to support flexible knowledge source selection at each atomic operator, and construct BlendQA to evaluate cross-knowledge source reasoning.
@@
Retrieval-augmented LLMs have shown promising potential in knowledge-intensive reasoning. Early works adopt a one-time retrieval strategy (Izacard et al., 2023b; Borgeaud et al., 2022), where the retriever is only called once to solve each question. This is often insufficient for solving multi-hop questions, which requires retrieval at each reasoning step. Hence, more recent approaches turn to multi-time retrieval strategies. IRCoT (Trivedi et al., 2023) and ITER-RETGEN (Shao et al., 2023) iteratively calls retrieval at each CoT step based on the previous steps’ results. Self-Ask (Press et al., 2023) iteratively generates a chain of sub-questions and performs retrieval at each sub-question. ProbTree (Cao et al., 2023) proposes to first formulate a global question decomposition plan in tree structure, then recursively perform retrieval at each sub-question. However, existing multi-time retrieval frameworks leverage LLMs to generate sub-questions in a free-form manner, imposing no supervision over each sub-question’s atomicity and functionality. Compared to previous methods, AtomR uses the combination of three predefined atomic knowledge operators to model complex reasoning procedures, achieving question decomposition that is highly fine-grained and orthogonal.
@@ 3. Methodology
In this section, we introduce the design of AtomR. The overall architecture of AtomR is illustrated in Figure 2. At the core of our framework are three fundamental Atomic Knowledge Operators (Section 3.1), which we design to retrieve and manipulate knowledge at the atomic level. These operators then steer the process of AtomR through two main stages: (1) Atomic Reasoning Planning (Section 3.2), where the system decomposes the input question into a fine-grained Atomic Reasoning Tree (ART), and (2) Atomic Reasoning Execution (Section 3.3), where the system performs bottom-up reasoning across multiple heterogeneous knowledge sources at each atomic node. We first introduce the design of our atomic knowledge operators in Section 3.1, then detail the two main stages of our framework in Sections 3.2 and 3.3. The LLM prompts of AtomR are included in Appendix H. The benchmark construction of BlendQA will be introduced later in Section 4.1.2.
@@ 3.1. Atomic Knowledge Operators
Compositional reasoning, the ability to construct complex reasoning procedures by combining simple components, is a fundamental property of human thought (Fodor and Pylyshyn, 1988; Fodor and Lepore, 2002). Knowledge graphs (KGs) effectively model the compositionality of knowledge, breaking down complex world knowledge into combinations of atomic elements: entities, relations, and attributes. Furthermore, KG query languages use predefined operations to manipulate the atomic elements in KGs. In this light, we ask the question: Can we use a set of predefined operators to atomically manipulate knowledge and enhance LLM compositional reasoning?
We investigate the essential operations of existing KG query languages, and observe that all operations can be distilled into combinations of three basic atomic knowledge operators: Search, Relate, and Filter. Table 1 illustrates how operations of three widely-used KG query languages, SPARQL, Cypher, and KoPL (Pérez et al., 2009; Francis et al., 2018; Cao et al., 2022), can be induced into AtomR’s three atomic knowledge operators. Each operator corresponds to a unique, orthogonal function. We implement each operator on multiple heterogeneous knowledge sources with an adaptive LLM executor, as shown in Figure 2(2)(iii).
@@
During execution, similar as Search, Relate first initiates multi-source knowledge retrieval using the concatenated parameters as the query, then inputs the retrieved knowledge and corresponding sub-question into the adaptive LLM executor with an in-context learning prompt. Finally, the LLM executor outputs the answer: either an entity list for case (1), an attribute value for case (2), or a relation for case (3).
@@
The overlapping coefficient of each entity is then compared to a hyperparameter threshold $t$. Entities with $O_{i}<t$ will be directly discarded to remove excessive noise from the retrieved knowledge. Then, the remaining entities and passages are inputted to the adaptive LLM executor to perform entity filtering via in-context learning. Finally, the LLM executor outputs the filtered entity list.
@@ 3.3.1. Leaf Node Reasoning
Finally, with the retrieved knowledge, AtomR executes the atomic knowledge operator. The detailed functionality and procedures of each atomic knowledge operator has been explained in Section 3.1. To ensure flexibility and robustness, we utilize an Adaptive LLM Executor, instead of static symbolic code, for atomic operator execution. Specifically, we employ in-context learning to prompt an LLM to perform the desired knowledge operation (i.e., Figure 8). Finally, the output of the adaptive LLM executor serves as the local answer for the atomic operator as well as its associated sub-question.
@@ 3.3.2. Parent Node Reasoning
[…]
where $q_{i}$ is the sub-question for the current node and $q_{c_{1}},a_{c_{1}},$...] is the list of question-answer pairs for $q_{i}$’s child nodes. For example, in Figure 2(1), sub-question 1 can be answered by synthesizing the answers of its child nodes 4 and 5. Child answer reasoning is achieved through in-context learning, where the LLM is provided with the child question-answer pairs along with an instruction.
@@
[…]
where $q_{s_{1}},a_{s_{1}},$...] is the list of question-answer pairs for $q_{i}$’s sibling nodes that are referenced in the current question $q_{i}$. For example, in Figure 2(1), sub-question 3 can be answered by analyzing answers of its referenced sibling node 2. Sibling answer reasoning is also achieved through LLM in-context learning.
@@
Each parent node is primarily resolved through either child answer reasoning or sibling answer reasoning. However, if a parent node fails to obtain an answer, direct RAG reasoning is employed. Parallel to leaf node reasoning, direct rag reasoning first performs dynamic knowledge source selection for the current sub-question $q_{i}$, then initiates multi-source knowledge retrieval with $q_{i}$ as the query. Finally, given $q_{i}$ and the retrieved knowledge, an LLM is employed to conduct standard RAG through in-context learning. This design ensures that while AtomR primarily focuses on knowledge retrieval at the leaf nodes, it can also flexibly access external knowledge at inner nodes when necessary.
@@ 4. Experiments
Table 2. Results for Single-Source Reasoning. Token-level F1 is reported, with the best results in bold and the second best results underlined. $(\uparrow)$ indicates the overall gain of AtomR compared to the second-best baseline for each dataset.
@@
AtomR (ours) 71.27(5.4)($\uparrow$ 5.4) 68.96 82.07 78.72(9.4)($\uparrow$ 9.4) 59.07 80.04 97.48 93.33 36.11(1.2)($\uparrow$ 1.2) 45.63 29.94 20.15
@@
AtomR (ours) 43.32(9.5)($\uparrow$ 9.5) 45.63 39.69 45.22 71.39(6.6)($\uparrow$ 6.6) 70.10 68.51 81.71 64.95